# Experimental Evaluation of a Two-Way Intelligibility Protocol for Human-LLM Interaction

### v2

@@ 6 Conclusion
Recently, we have seen a dramatic expansion in the use of ML, enabling it to support almost any activity where data can be collected and analysed. A challenge arises when the predictive power of modern ML meets the need for human understanding. On the face of it, the use of machine-agents that communicate in natural language can mitigate the problem by providing explanations in a human-readable form. But, readability, while necessary for understandability, may not be sufficient. In this paper, we investigate the use of an information-exchange protocol specifically designed around the notion of intelligibility of messages exchanged between agents, that attempts to address the broader issue of the quality of information exchanged in human–machine communication Coiera (2001).
The concept of “intelligibility of communication” underpins the abstract interaction model proposed in Baskar et al. (2024), which explores how human–ML systems can collaborate to predict and explain data. While that work is conceptual, with no implementation, it presents case studies showing how an “intelligibility protocol” could qualitatively assess communication clarity between humans and machines.
This paper builds on that idea by implementing a simple interactive procedure that exchanges messages as proposed in Baskar et al. (2024) and testing it on problems involving human interaction with large language models (LLMs). LLMs are well-suited for this because (1) their natural language capabilities enable effective collaboration with non-ML experts who have domain expertise, and (2) foundational LLMs possess extensive general knowledge useful for complex, data-driven analysis. The results are a first step toward empirical support for using intelligibility as a foundation for designing collaborative human–ML systems.
@@ Appendix A Experiment Details
All experiments use claude-3-5-sonnet-latest Anthropic (2024) as the LLM. For ASK_AGENT (Algorithm 2). The max_tokens for RAD is set to 300, and for DRUG it is set to 1024. For AGREE, the temperature parameter for both RAD and DRUG is set to 0, and max_tokens is set to 10. All other settings are the defaults, set by the Claude API.
@@ Appendix B Dataset Details
RAD: We use Radiopaedia  Radiopaedia (2007) for data. This is a multi-modal dataset compiled and peer-reviewed by radiologists. Each tuple in our database consists of: (a) An X-ray image; (b) A prediction consisting of a set of diagnosed diseases, and (c) an explanation in the form of a radiological report. Both (b) and (c) are from expert human annotators. We use a subset of the full Radiopedia database, focusing on 5 ailments ($Atelectasis$, PneumoniaPneumonia, PneumothoraxPneumothorax, PleuralEffusionPleuralEffusion, and $Cardiomegaly$) as our database with 4 instances per disease (20 instances overall). We summarise the reports using gpt-3.5-turbo-0125 and use these as the ground truth.
DRUG: Molecules are selected from DrugBank Wishart et al. (2018) (drugs are leads that satisfy additional biological and commercial constraints). It contains a list of drug molecules approved by the FDA (Food and Drug Administration). Specifically, we will focus on a subset of molecules with a molecular weight between 150g/mol - 300g/mol and have at least one aromatic ring. This set has been further sampled to get 20 molecules.

### v4

@@ 6 Conclusion
Recently, we have seen a dramatic expansion in the use of ML, enabling it to support almost any activity where data can be collected and analysed. A challenge arises when the predictive power of modern ML meets the need for human understanding. On the face of it, the use of machine-agents that communicate in natural language can mitigate the problem by providing explanations in a human-readable form. But, readability, while necessary for understandability, may not be sufficient. In this paper, we investigate the use of an information-exchange protocol specifically designed around the notion of intelligibility of messages exchanged between agents, that attempts to address the broader issue of the quality of information exchanged in human–machine communication [5].
The concept of “intelligibility of communication” underpins the abstract interaction model proposed in [3], which explores how human–ML systems can collaborate to predict and explain data. While that work is conceptual, with no implementation, it presents case studies showing how an “intelligibility protocol” could qualitatively assess communication clarity between humans and machines.
This paper builds on that idea by implementing a simple interactive procedure that exchanges messages as proposed in [3] and testing it on problems involving human interaction with large language models (LLMs). LLMs are well-suited for this because (1) their natural language capabilities enable effective collaboration with non-ML experts who have domain expertise, and (2) foundational LLMs possess extensive general knowledge useful for complex, data-driven analysis. The results are a first step toward empirical support for using intelligibility as a foundation for designing collaborative human–ML systems.
@@ Appendix A Experiment Details
All experiments use claude-3-5-sonnet-latest [2] as the LLM. For ASK_AGENT (Algorithm 2). The max_tokens for RAD is set to 300, and for DRUG it is set to 1024. For AGREE, the temperature parameter for both RAD and DRUG is set to 0, and max_tokens is set to 10. All other settings are the defaults, set by the Claude API.
@@ Appendix B Dataset Details
RAD: We use Radiopaedia  [13] for data. This is a multi-modal dataset compiled and peer-reviewed by radiologists. Each tuple in our database consists of: (a) An X-ray image; (b) A prediction consisting of a set of diagnosed diseases, and (c) an explanation in the form of a radiological report. Both (b) and (c) are from expert human annotators. We use a subset of the full Radiopedia database, focusing on 5 ailments ($Atelectasis$, PneumoniaPneumonia, PneumothoraxPneumothorax, PleuralEffusionPleuralEffusion, and $Cardiomegaly$) as our database with 4 instances per disease (20 instances overall). We summarise the reports using gpt-3.5-turbo-0125 and use these as the ground truth.
DRUG: Molecules are selected from DrugBank [18] (drugs are leads that satisfy additional biological and commercial constraints). It contains a list of drug molecules approved by the FDA (Food and Drug Administration). Specifically, we will focus on a subset of molecules with a molecular weight between 150g/mol - 300g/mol and have at least one aromatic ring. This set has been further sampled to get 20 molecules.